Consider a series of `n` concentric circles C(1),C(2),C(3), ....., C(n) with radii r(1),r(2),r(3), ......., r(n) respectively, satisfying r(1)>r(2)>r(3).... > r(n) & r(1)=10. The circles are such that the chord of contact of tangents from any point on Ci to C(i+1) is a tangent to C(i+2) (i=1,2,3,...). Find the value of r(1)+r(2)+r(3)+.........r(infinity), if the angle between the tangents from any point of C(1) to C(2) is 60.

Grade:7

3 Answers

Arun Kumar IIT Delhi
9 years ago
Hi
Thanks & Regards, Arun Kumar, Btech,IIT Delhi, Askiitians Faculty
Arun Kumar IIT Delhi
9 years ago
Hi
consider making the diagram yourself
you must know that c(n) is the innermost circle
i'm assuming tangents form c(i) to c(i+1) of which chord
of contact touches c(i+2)
try to understand the analysis
consider the triangle made by r(i+1),r(i) and length of tangent
angle between length of tangent and r(i)=30
(since 60 is given)
so sin30={r(i+1) / r(i)}
r(i+1)={r(i) / 2}
consider one more triangle by r(i+1),half of chord and r(i+2)
again cos(60)={r(i+2) /r(i+1)}
=>r(i+2)={r(i+1) / 2}
=>GP with common ratio 1/2
=>sum={a / 1-r}={10 / (1-1/2)}=20
Thanks & Regards, Arun Kumar, Btech,IIT Delhi, Askiitians Faculty
Arun Kumar IIT Delhi
9 years ago
Hello
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty